An insurance company offers Mississippi adults between the ages of 25 and 34 a $100,000 life insurance policy for $18 a month. They use the fact that Mississippi has a yearly death rate of 165 per 100,000 residents aged 25 – 34 years.

Find the expected value per customer for the insurance company at the end of one year for the policy described.

If the insurance company has 5,000 customers with these life insurance policies in Mississippi, what is its profit at the end of the year
Answer #1

Answer: Given,

p (death) = 0.00165

E(X) for 1 year = ?

Since it is for 1 year, we will have 12 installments of $18. So:

= p1*x1 +p2*x2

= 0.00165*100000 - (1-0.00165)*18*12

= 165 - 215.6436

=$50.6436

So, the expected value is a profit of of $50.6436 per customers for the insurance company at the end of 1 year of the stated policy.

Note: As per the HomeworkLib policy i can answered only 1 as early as possible i will update another question .